NPP Projection Admin 2/3
To do this evening/ was completed: LMS scheduled for coming week with some first run exceptions.
Tech'd Frankenhooker, Removal of the Eye, Monk in Pieces, Let the Right One In. Volumes logged, playlists updated. Pillion and GOAT ingested to LMS.
Loops ran, projectors cleaned.
Pulp Fiction trailer back on its core. Die Hard packaged up. Asked John to double check return shipping location. Second can of the SoF print arrived, it's in T1.
*A1 2/6 playlists need to be updated with the Mubi La Grazia trailer sometime in the next few days.*
T6 storage is at maximum capacity, and there isn't much I can remove right now. We are fine until Monday 2/9 when something will have to be removed to make room for Sinners.
Attempted to download the Dead Lover DCP but the Aspera link was prompting me to commit to some new transfer client download so I held off for the moment.
Presentation Issues:
Preshow Clips/Playlist:
Trailer Updates:
Special Events/ AV set up:
Ordering/Receiving (posters & trailers):
Ordering/Receiving (content):
Outgoing/Shipping (content):
Print/video Inspection:
Asset tracking doc updated/checked?:
Maintenance/cleaning:
Cue sheets updated/checked (every shift): Cue sheets for next week are done except for the litany of new titles that will be coming at us next Thursday since we don't have the timing info on any of those.
Marquee updated?:
Posters updated?:
Cleaning lights on for the night?:
Projection cart by entrance?:
Additional Notes:
Comments
Post a Comment